OpenStudy (precal):

Determine the open intervals on which the given function is increasing or decreasing and the coordinates of any relative extrema.

OpenStudy (precal):

\[f(x)=x ^{2/3}-4\]

OpenStudy (precal):

I took the derivative and set it equal to zero, but when I solved it I got 0 not equal to 2 not sure what I am doing incorrect

OpenStudy (precal):

\[f' (x)=\frac{ 2 }{ 3\sqrt[3]{x} }\]

zepdrix (zepdrix):

0 not equal to 2? I'm not sure what that means :o

OpenStudy (precal):

\[\frac{ 2 }{ 3\sqrt[3]{x} }=0\]

zepdrix (zepdrix):

Your derivative looks good. Looking for critical points,\[\Large\rm 0=\frac{2}{3}x^{-1/3}\]Multiply each side by 3/2,\[\Large\rm 0=x^{-1/3}\]So it looks like we get a critical point at x=0, yes? :o

OpenStudy (precal):

boy do I feel dumb today

OpenStudy (precal):

now I know how to finish this problem can you help me solve one more?

zepdrix (zepdrix):

sure c:

OpenStudy (precal):

\[h(x)=\frac{ x^2-3x-4 }{ x-2 }\]

OpenStudy (precal):

same directions so I took the derivative and got

OpenStudy (precal):

\[h'(x)=\frac{ x^2-4x+10 }{ (x-2)^2 }\]

OpenStudy (precal):

so is my critical point x=2 or I thought I set it equal to zero and solve for x? any case that meant I have to solve the numerator and I got a complex solution so what am I overlooking again?

zepdrix (zepdrix):

We deal with the numerator and denominator separately. The denominator is giving us critical points for a different reason than we're used to. It's not giving us locations of horizontal slope. This was true for the last problem as well, the critical point ended up being a cusp. The problem works out a little differently, you find x=2 to be a critical point. But wait! We have to ask ourselves, "Is x=2 in the DOMAIN of my function?"

OpenStudy (precal):

so the last problem does not have a relative extrema because x=0 is not differentiable. and you are right x =2 is not in the domain

zepdrix (zepdrix):

does not have relative extrema because the function is not differentiable at x=0, Ooo yah I like the way you said that! :O The last one `did` have a critical point, but not the kind we were hoping for >.<

OpenStudy (precal):

so the critical point is complex? or in this case undefined

OpenStudy (precal):

but what I am to write, I am lost on this one

zepdrix (zepdrix):

For the second one? No critical points exist within the domain of the function, therefore no extreme points exist either. Or however you think that should be worded. Oh oh oh, but you still need to determine increasing/decreasing, yes?

OpenStudy (precal):

yes, should I test point on both sides of 2 since I know that x=2 is the vertical asymptote and is an undefined point?

zepdrix (zepdrix):

Yah that seems like a good idea >.<

OpenStudy (precal):

thanks, I couldn't think of anything else.....thank you so much

OpenStudy (precal):

only 500 more impossible calculus problems to go.....

zepdrix (zepdrix):

Try to simplify things if you can. Since we only care about the `sign` of the result, notice that the denominator is always positive, so it's not telling you anything:\[\Large\rm h'(x)=\frac{ x^2-4x+10 }{ positive }\]And then plug the `less than 2` or `greater than 2` value into the top to see what's going on.

zepdrix (zepdrix):

Impossible?? You mean super funnnn? \c:/

OpenStudy (precal):

yes, I try to do them as best I can and I come here to double check my solutions, thoughts, process, etc........